# HyConEx: Hypernetwork classifier with counterfactual explanations

### v1

@@ 4.4 Ablation Study
[…]
Base 1.00 $0$
Base+CE 1.00 1.000 6.394 1.628 $0$34761532.00 3.925 -0.036
Base+CE+Flow 1.00 1.000 1.461 0.475 0.662 33.707 1.108 0.058
Base+CE+Dist 1.00 0.988 0.098 0.032 0.394 27.66 1.072 0.060
Full 1.00 0.997 0.416 0.142 0.602 34.61 1.073 0.065
@@ 5 Conclusion
Understanding model decisions is crucial in explainable AI, mainly through counterfactual explanations highlighting key feature modifications affecting predictions. While existing methods generate counterfactuals for external models, no interpretable classifier has produced predictions and counterfactuals simultaneously. We introduced HyConEx, a deep hypernetwork-based model tailored for tabular data, which provides accurate classifications and generates counterfactual examples within high-density class regions. Our approach achieves competitive performance while ensuring interpretability, making HyConEx a unique all-in-one neural network that seamlessly integrates prediction and explanation.

### v2

@@ 5.4 Ablation Study
[…]
Base 1.00 0.000 - - - - - -
Base+CE 1.00 1.000 6.394 1.628 0.000 -34761532.00 3.925 -0.036
Base+CE+Flow 1.00 1.000 1.461 0.475 0.662 33.707 1.108 0.058
Base+CE+Dist 1.00 0.988 0.098 0.032 0.394 27.66 1.072 0.060
Full 1.00 0.997 0.416 0.142 0.602 34.61 1.073 0.065
@@ 6 Conclusion
Understanding model decisions is crucial in explainable AI, mainly through counterfactual explanations highlighting key feature modifications affecting predictions. While existing methods generate counterfactuals for external models, no interpretable classifier has produced predictions and counterfactuals simultaneously. We introduced HyConEx, a deep hypernetwork-based model tailored for tabular data, which provides reasonable classifications and generates counterfactual examples within high-density class regions. Our approach achieves competitive performance while ensuring interpretability, making HyConEx a unique all-in-one neural network that seamlessly integrates prediction and explanation.
Limitations
HyConEx is intentionally tailored for tabular datasets. This specialization yields strong performance where interpretability is most critical, but broadening the method to other data modalities would require additional architectural modifications. Although inference is highly efficient, this capability depends on an upfront training phase; once completed, however, HyConEx supports real-time deployment and substantially outperforms prior methods in runtime.
HyConEx delivers a single counterfactual candidate for every alternative class. While this approach allows us to return a plausible counterfactual example with good proximity, we do not provide a user with multiple candidates as in the case of generative models. Moreover, since counterfactual explanations are embedded into HyConEx classifier, we do not take into account the change of the classification decision boundary. This property refers to the ability of counterfactual examples to withstand small perturbations in the classifier. Additional modifications of HyConEx would be required to produce counterfactuals, which are robust to such modifications.
Future work
The concept of a hypernetwork generating local linear decision boundaries, introduced by IMN and further developed in HyConEx, demonstrates the considerable potential of this type of architecture. Therefore, future work on this framework can progress through both incremental refinements and more transformative innovations.
In particular, it is worth to consider different hypernetwork backbones, such as transformer-based (e.g., FT-Transformer [6]), retrieval-augmented (TabR [41]) or ensemble-like (TabM [42]) models. Additionally, we could potentially try using other differentiable density estimation models, especially for high-dimensional or categorical datasets. Finally, it would be valuable to extended the scope of HyConEx to other modalities, such as images or texts.
Acknowledgment
The research of P. Marszałek, U. Movsum-zada supported by the National Science Centre (Poland), grant no. 2023/50/E/ST6/00169. The work of K. Książek was supported by the flagship project ”Artificial Intelligence Computing Center Core Facility” from the DigiWorld Priority Research Area within the Excellence Initiative – Research University program at Jagiellonian University in Krakow. The work of P. Spurek was supported by the National Science Centre (Poland), Grant No. 2023/50/E/ST6/00068. The work of M. Śmieja was supported by the National Science Centre (Poland), Grant No. 2022/45/B/ST6/01117. Oleksii Furman work was supported by the National Science Centre (Poland) Grant No. 2024/55/B/ST6/02100.